OpenStudy (anonymous):

Trevor's dad decided to get him a cell phone for his birthday. Cell Plus has a plan that costs $27.00 per month plus an additional $0.06 per minute. Cell Best has a plan that costs $30.00 per month plus an additional $0.05 per minute. How many minutes can Trevor talk and both cell plans cost the same amount? 600 30 150 300

OpenStudy (anonymous):

Let x be minutes that Trevor spends phoning Cell plus costs: \[27+0.06x\]Whereas Cell best costs \[30+0.05x\]If we set those two as equal\[27+0.06x=30+0.05x\]\[0.01x=3\]\[x=300\]

OpenStudy (anonymous):

\[0.06min + 27.00 = 0.05min + 30.00\] \[0.01min = 3.00\] \[min = 300 \text{ minutes}\]